# Near–Real-Time Conflict-Related Fire Detection Using Unsupervised Deep Learning and Satellite Imagery

### v2

@@ Abstract
Ongoing armed conflict in Sudan highlights the need for rapid monitoring of conflict-related fire damage. Recent advances in deep learning and high-frequency satellite imagery enable near–real-time assessment of active fires and burn scars in war zones. This study presents a near–real-time monitoring approach using a lightweight Variational Auto-Encoder (VAE)–based model integrated with 4-band Planet Labs imagery at 3 m spatial resolution. We demonstrate that conflict-related fire damage can be detected with minimal delay using accessible, commercially available satellite data. To achieve this, we adapt a VAE-based model, originally designed for 10-band imagery, to operate effectively on high-resolution 4-band inputs. The model is trained in a unsupervised manner to learn compact latent representations of nominal land-surface conditions and identify fire-affected areas by quantifying changes between temporally paired latent embeddings. Performance is evaluated across five case studies in Sudan and compared against a cosine-distance baseline computed between temporally paired image tiles using precision, recall, F1-score, and the area under the precision–recall curve (AUPRC). Results show that the proposed approach consistently outperforms the baseline, achieving higher recall and F1-scores while maintaining strong precision in highly imbalanced fire-detection scenarios. Experiments with 8-band imagery and temporal image sequences yield only marginal performance gains over single 4-band inputs, underscoring the effectiveness of the proposed lightweight approach for scalable, near–real-time conflict monitoring.
@@
journal: Remote Sensing Applications: Society and Environment
@@ 1 Introduction
An armed conflict in Sudan, which began in April 2023, has resulted in widespread civilian harm, large-scale displacement, and severe destruction of infrastructure. Fighting between the Sudanese Armed Forces (SAF) and the Rapid Support Forces (RSF) has resulted in the deaths of thousands of people and the displacement of over 12 million individuals (Birch et al., 2024). The conflict initially concentrated in Khartoum state before expanding to western and southern regions, particularly Darfur, where attacks on civilians, healthcare facilities, and essential infrastructure are extensively documented (Milton et al., 2025; Dahab et al., 2025). (Eljack et al., 2023) and (Alrawa et al., 2023) highlight the impact of conflict on health.
The Sudan Conflict Observatory (SCO) has reported a massacre of more than a thousand civilians by June 2023 in El-Geneina (Rothbart et al., 2025), and reported destruction of healthcare facilities (Abubakr et al., 2024). Figure 1 shows the map of attacks in Sudan. The timeline indicates a rising number of attacks since April 2023, with a significant increase in the regions of Khartoum, El Fasher, and El Geneina.
Figure 1: Map of attacks in Sudan, April–June 2023, highlighting the key conflict hotspots.
@@ 1.1 Problem Statement and Objectives
Monitoring conflict-related fires in active war zones presents several challenges. Ground-based reporting is often delayed, incomplete, or impossible due to insecurity and access constraints (Sticher et al., 2023). As a result, satellite imagery has become a critical source of independent evidence for tracking attacks, infrastructure destruction, and potential violations of international humanitarian law (Hassan and Ahmed, 2025). Among observable conflict evidence, active fires and burn scars are particularly informative, as they frequently accompany airstrikes, shelling, looting, and the destruction of civilian structures. Recent advances in high-resolution, high-frequency satellite imagery, combined with deep learning techniques, enable the possibility of near–real-time conflict monitoring. Commercial satellite constellations, such as Planet Labs, provide near-daily revisit capabilities at meter-scale resolution, which makes them particularly suitable for detecting small-scale fires in dense environments. Similarly, deep learning methods provide a way to detect anomalous events without relying on ground-truth labels, which are limited in active conflict zones (Aung et al., 2021). However, existing fire-detection and damage-assessment methods face key limitations in conflict settings. Many approaches rely on coarse-resolution sensors, supervised learning pipelines, or computationally intensive models that are not easily scalable for rapid deployment (Racek et al., 2025). Moreover, models trained on wildfire datasets often perform poorly in urban conflict zones, where fires are short-lived, spatially fragmented (Mueller et al., 2021; Di Martino et al., 2023), and embedded in heterogeneous landscapes (Pleniou and Koutsias, 2025; Franquesa et al., 2022). These constraints hinder timely detection of fire-affected areas, which is critical for humanitarian assessment and operational decision-making.
@@
    What is the marginal utility of short temporal sequences and additional spectral bands compared to single-snapshot multispectral imagery?
The scope of this study is limited to satellite-based observations in Sudan and focuses on the detection of fires and burn scars. Retrospective damage assessment and causal attribution are beyond the scope of this work. The framework is designed for near–real-time monitoring, prioritizing the recall of fire-affected areas while maintaining operationally viable precision. We assume the availability of cloud-free imagery or the application of standard cloud-masking techniques, recognizing that dense cloud cover remains a limitation of optical remote sensing.
The specific objectives of this study are to:
@@
Collectively, these contributions provide an operationally viable, unsupervised approach for near–real-time detection of conflict-related fire-affected areas. Furthermore, the study empirically quantifies the marginal benefits of short temporal sequences and additional spectral bands for improving detection performance in conflict settings.
@@ 2.1 Deep Learning in Fire Detection
Convolutional Neural Networks (CNNs) are widely adopted for automating fire detection from satellite imagery by identifying thermal anomalies or spectral signatures of burn scars (Toan et al., 2019; Seydi et al., 2022). Architectures like U-Net improve pixel-wise segmentation, enabling finer distinctions between active fires and background vegetation (Hally et al., 2019). More recently, Vision Transformers (ViTs) have demonstrated improved performance in capturing long-range spatial dependencies compared to traditional CNNs (Saleh et al., 2024). However, these supervised models remain heavily dependent on high-quality, labeled training data, which is scarce in active conflict zones. Moreover, while medium-resolution sensors like Moderate Resolution Imaging Spectroradiometer (MODIS) and Visible Infrared Imaging Radiometer Suite (VIIRS) are effective at thermal sensing, they often struggle to capture small-scale fires in dense or irregular settlements (Bromley, 2010; Zhao et al., 2023).
@@ 2.2 Challenges in Conflict Monitoring and Damage Assessment
Automated satellite-based conflict monitoring has advanced substantially, yet manual interpretation remains prevalent for assessing conflict-related damage (Avtar et al., 2021). Early studies relied on handcrafted features derived from Very High Resolution (VHR) imagery for post-conflict damage assessment, such as in the Gaza Strip (Kahraman et al., 2016). Subsequent studies applied supervised and semi-supervised learning approaches to detect damaged infrastructure in Syria using VHR imagery (Lee et al., 2020; Mueller et al., 2021). While effective, these methods depend on expensive imagery, labeled training data, and retrospective pre- and post-event comparisons, limiting their scalability and timeliness over large geographic extents. Research in Sudan and Darfur has historically utilized MODIS data, which, despite its efficacy in detecting large-scale events, struggles to identify the smaller, dispersed fires typical of irregular warfare (Bromley, 2010; Cornebise et al., 2018). Time-series and change detection methods have also been explored in conflict contexts (Jenerowicz et al., 2010; Braun, 2018). However, these approaches typically focus on post-conflict or long-term damage assessment rather than near–real-time fire detection, highlighting the need for scalable, unsupervised methods capable of rapid detection, as proposed in this study. Case studies across conflict and disaster settings, including Myanmar (Marx et al., 2019) and Turkey (Kurnaz et al., 2020), demonstrate the potential of satellite imagery for fire detection, yet a gap remains in fully unsupervised methods capable of processing high-resolution data streams without labels (Xu et al., 2022).
@@ 2.3 Unsupervised Learning and Anomaly Detection
To mitigate label dependency, recent work has increasingly explored unsupervised anomaly detection, in which models learn the distribution of nominal land cover and identify deviations without requiring explicit annotations (An and Cho, 2015). Variational Autoencoders (VAEs) are frequently employed in this setting due to their ability to learn compact latent representations of complex imagery (Kingma et al., 2013). While many approaches rely on reconstruction error as an anomaly signal, prior studies have shown that reconstruction-based scores can be unreliable in fully unsupervised settings, particularly for high-capacity models (Merrill and Eskandarian, 2020). More recent work has therefore emphasized the use of latent representations for comparing imagery across time, framing anomaly detection as a form of change detection in latent space (Bergamasco et al., 2022). This approach is particularly relevant in conflict zones, where unprecedented events and limited historical labels make representation-based change detection more robust than reconstruction-based anomaly scoring.
@@ 2.4 Sensor Selection and Data Processing
Satellite sensor choice determines the trade-off between spatial resolution and temporal frequency. Landsat and Sentinel-2 provide valuable multispectral data for burn scar analysis; however, their revisit intervals of 5–10 days often miss short-lived fire events (Schroeder et al., 2016). Synthetic Aperture Radar (SAR) can penetrate cloud cover to assess post-fire land changes, though it lacks the spectral signatures provided by optical sensors (Narvaez Luces et al., 2023). PlanetScope imagery offers a balance, providing 3 m resolution and near-daily revisit cadences. However, leveraging this data for conflict monitoring requires robust preprocessing and computationally efficient models capable of handling large daily data volumes, factors that have limited its adoption in fully unsupervised, near–real-time fire detection workflows (Boroujeni et al., 2024; de Almeida Pereira et al., 2021; Bazi et al., 2021).
@@ 3 Study area
This study focuses on five conflict-related fire incidents in the El Fasher and Khartoum regions of Sudan. In Khartoum State, an airstrike reportedly carried out by the Sudanese Armed Forces (SAF) on May 25, 2024, targeted the Gandahar Market in Omdurman. The Conflict Observatory documented damage to at least 31 structures resulting from this incident (Etefa, 2026). PlanetScope imagery acquired on the day of the strike shows a visible smoke plume over the market area, consistent with active burning.
In North Darfur, political and security tensions escalated in late March 2024, with both the SAF and the Rapid Support Forces (RSF) preparing for confrontation in El Fasher by early May. According to the Conflict Observatory, at least 7,800 structures in eastern El Fasher were affected by fighting beginning on May 9, 2024, with NASA FIRMS data indicating widespread fire activity and extensive burning by May 11 (Rothbart et al., 2024a). Earlier, in April 2024, RSF-led forces attacked approximately 17 villages surrounding El Fasher. These attacks caused significant civilian casualties and widespread destruction of residential structures (Sabahelzain et al., 2025). PlanetScope imagery captured active fires in several of these locations on April 13, 2024, including the villages of Jaranga, Muqrin, and Sarafaya (Rothbart et al., 2024b).
Table 1 summarizes the imagery and incident details for all case studies considered in this work. Additionally, Figure 2 illustrates the geographic distribution of the study areas across Sudan. These incidents were selected based on independently reported fire-related damage documented in (Howarth et al., 2024a; Sabahelzain et al., 2025; Howarth et al., 2024b), and manual evaluation and verification conducted by analysts associated with the Sudan Conflict Observatory project.
@@ 4.1 Model architecture
In this study, we adapted and retrained RaVAEn (Růžička et al., 2022), a lightweight convolutional Variational Autoencoder (VAE) framework originally developed for unsupervised detection of disaster-related surface anomalies across multiple hazard types, including floods, fire burn scars, landslides, and hurricanes. Rather than being tailored to a specific phenomenon, RaVAEn learns a compact latent representation of nominal surface conditions and identifies anomalous patterns as deviations from this learned distribution, making it suitable for detecting previously unseen or irregular events such as conflict-related fires. The original model was designed for 10-band satellite imagery at 10 m spatial resolution. We retrained it for 3 m resolution imagery with four spectral channels (RGB and near-infrared) to better capture fine-scale spatial heterogeneity associated with small, fragmented fire events in urban conflict settings.
@@
To accommodate scale-variant spatial patterns, standard strided convolutions in the initial stages are replaced with multi-scale convolutional blocks analogous to Atrous Spatial Pyramid structures (Maggiori et al., 2016). These blocks use parallel $3\times 3$ convolutions with increasing dilation rates, fused via a $1\times 1$ convolution. This design enables the encoder to capture multi-scale contextual dependencies and expand its effective receptive field while preserving spatial detail (Chen et al., 2017; Yuan et al., 2018). Anti-aliased downsampling, or BlurPool (Zhang, 2019), implemented as a stride-1 convolution followed by low-pass filtering and pooling, mitigates the loss of high-frequency information and improves shift equivariance when processing higher-resolution imagery. Each downsampling operation is followed by residual convolutional blocks with leaky ReLU activations and normalization layers, while deeper stages retain a conventional residual structure as higher-level semantic features are less sensitive to pixel-level resolution.
The final encoder feature map is aggregated using global pooling and projected onto a 128-dimensional latent space parameterized by a diagonal Gaussian distribution, regularized with a $\beta$-VAE objective ($\beta$=1). The decoder mirrors the encoder hierarchy using nearest-neighbor upsampling and convolutional layers, with a linear activation in the output layer. The model is trained end-to-end using the Adam optimizer with a learning rate of 0.001 and a batch size of 128, minimizing a combination of mean squared reconstruction loss and Kullback–Leibler divergence. To improve generalization across spatial resolutions, a scale-augmented training strategy is applied in which imagery is stochastically resampled to simulate continuous transitions in ground sampling distance (Koziarski and Cyganek, 2017). This resampling is applied only during training; inference operates directly on native-resolution imagery, producing resolution-invariant feature embeddings.
While VAEs are commonly used for unsupervised anomaly detection by thresholding reconstruction error for individual inputs (Angerhausen et al., 2022), prior work has shown that reconstruction error can be an unreliable indicator of anomalous behavior in fully unsupervised settings (Merrill and Eskandarian, 2020). Motivated by these findings, we do not use reconstruction error as the anomaly signal. Instead, the VAE is used as a representation learning framework to embed surface conditions into a compact latent space, in which changes between temporally paired observations can be quantified for change detection.

### v3

# arXiv:2512.07925v3[cs.CV] 01 Mar 2026

@@ Abstract
Ongoing armed conflict in Sudan highlights the need for rapid monitoring of conflict-related fire-affected areas. Recent advances in deep learning and high-frequency satellite imagery enable near–real-time assessment of active fires and burn scars in war zones. This study presents a near–real-time monitoring approach using a lightweight Variational Auto-Encoder (VAE)–based model integrated with 4-band Planet Labs imagery at 3 m spatial resolution. We demonstrate that these impacted regions can be detected within approximately 24 to 30 hours under favorable observational conditions using accessible, commercially available satellite data. To achieve this, we adapt a VAE-based model, originally designed for 10-band imagery, to operate effectively on high-resolution 4-band inputs. The model is trained in an unsupervised manner to learn compact latent representations of nominal land-surface conditions and identify burn signatures by quantifying changes between temporally paired latent embeddings. Performance is evaluated across five case studies in Sudan and compared against cosine distance, CVA, and IR-MAD using precision, recall, F1-score, and the area under the precision–recall curve (AUPRC) computed between temporally paired image tiles. Results show that the proposed approach consistently outperforms the other methods, achieving higher recall and F1-scores while maintaining operationally viable precision in highly imbalanced fire-detection scenarios. Experiments with 8-band imagery and temporal image sequences yield only marginal performance gains over single 4-band inputs, underscoring the effectiveness of the proposed lightweight approach for scalable, near–real-time conflict monitoring.
@@
journal: Science of Remote Sensing
@@ 1 Introduction
An armed conflict in Sudan, which began in April 2023, has resulted in widespread civilian harm, large-scale displacement, and severe destruction of infrastructure. Fighting between the Sudanese Armed Forces (SAF) and the Rapid Support Forces (RSF) has resulted in the deaths of thousands of people and the displacement of over 12 million individuals [11]. The conflict initially concentrated in Khartoum state before expanding to western and southern regions, particularly Darfur, where attacks on civilians, healthcare facilities, and essential infrastructure are extensively documented [47, 19]. The impact of conflict on health has been highlighted in recent studies [24, 3].
The Sudan Conflict Observatory (SCO) has reported a massacre of more than a thousand civilians by June 2023 in El-Geneina [55], and reported destruction of healthcare facilities [1]. Figure 1 shows the map of attacks in Sudan. The timeline indicates a rising number of attacks since April 2023, with a significant increase in the regions of Khartoum, El Fasher, and El Geneina.
Figure 1: Map of attacks in Sudan, April 2023–June 2024, highlighting the key conflict hotspots.
@@ 1.1 Problem Statement and Objectives
Monitoring conflict-related fires in active war zones presents several challenges. Ground-based reporting is often delayed, incomplete, or impossible due to insecurity and access constraints [61]. As a result, satellite imagery has become a critical source of independent evidence for tracking attacks, infrastructure destruction, and potential violations of international humanitarian law [31]. Among observable conflict evidence, active fires and burn scars are particularly informative, as they frequently accompany airstrikes, shelling, looting, and the destruction of civilian structures. Recent advances in high-resolution, high-frequency satellite imagery, combined with deep learning techniques, enable the possibility of near–real-time conflict monitoring. Commercial satellite constellations, such as Planet Labs, provide near-daily revisit capabilities at meter-scale resolution, which makes them particularly suitable for detecting small-scale fires in dense environments. Similarly, deep learning methods provide a way to detect anomalous events without relying on ground-truth labels, which are limited in active conflict zones [6]. However, existing fire-detection and damage-assessment methods face key limitations in conflict settings. Many approaches rely on coarse-resolution sensors, supervised learning pipelines, or computationally intensive models that are not easily scalable for rapid deployment [52]. Moreover, models trained on wildfire datasets often perform poorly in urban conflict zones, where fires are short-lived, spatially fragmented [48, 21], and embedded in heterogeneous landscapes [51, 27]. These constraints hinder timely detection of fire-affected areas, which is critical for humanitarian assessment and operational decision-making.
@@
    What is the marginal utility of short temporal sequences or additional spectral bands compared to single-snapshot multispectral imagery?
The scope of this study is limited to satellite-based observations in Sudan and focuses on the detection of fires and burn scars. Retrospective damage assessment and causal attribution are beyond the scope of this work. The framework is designed for near–real-time monitoring, prioritizing the recall of fire-affected areas while maintaining operationally viable precision. We assume the availability of cloud-free imagery or the application of standard cloud-masking techniques, recognizing that dense cloud cover remains a limitation of optical remote sensing. The specific objectives of this study are to:
@@
Collectively, these contributions demonstrate the feasibility of an operational, unsupervised approach for near–real-time detection of conflict-related burn signatures. Furthermore, the study empirically quantifies the marginal benefits of short temporal sequences and additional spectral bands for improving detection performance in conflict settings.
@@ 2.1 Deep Learning in Fire Detection
Convolutional Neural Networks (CNNs) are widely adopted for automating fire detection from satellite imagery by identifying thermal anomalies or spectral signatures of burn scars [62, 60]. Architectures like U-Net improve pixel-wise segmentation, enabling finer distinctions between active fires and background vegetation [30]. More recently, Vision Transformers (ViTs) have demonstrated improved performance in capturing long-range spatial dependencies compared to traditional CNNs [58]. However, these supervised models remain heavily dependent on high-quality, labeled training data, which is scarce in active conflict zones. Moreover, while medium-resolution sensors like Moderate Resolution Imaging Spectroradiometer (MODIS) and Visible Infrared Imaging Radiometer Suite (VIIRS) are effective at thermal sensing, they often struggle to capture small-scale fires in dense or irregular settlements [14, 66].
@@ 2.2 Challenges in Conflict Monitoring and Damage Assessment
Automated satellite-based conflict monitoring has advanced substantially, yet manual interpretation remains prevalent for assessing conflict-related damage [7]. Early studies relied on handcrafted features derived from Very High Resolution (VHR) imagery for post-conflict damage assessment, such as in the Gaza Strip [35]. Subsequent studies applied supervised and semi-supervised learning approaches to detect damaged infrastructure in Syria using VHR imagery [39, 48]. While effective, these methods depend on expensive imagery, labeled training data, and retrospective pre- and post-event comparisons, limiting their scalability and timeliness over large geographic extents. Research in Sudan and Darfur has historically utilized MODIS data, which, despite its efficacy in detecting large-scale events, struggles to identify the smaller, dispersed fires typical of irregular warfare [14, 18]. Time-series and change detection methods have also been explored in conflict contexts [34, 13]. However, these approaches typically focus on post-conflict or long-term damage assessment rather than near–real-time fire detection, highlighting the need for scalable, unsupervised methods capable of rapid detection, as proposed in this study. Case studies across conflict and disaster settings, including Myanmar [43] and Turkey [38], demonstrate the potential of satellite imagery for fire detection, yet a gap remains in fully unsupervised methods capable of processing high-resolution data streams without labels [63].
@@ 2.3 Unsupervised Learning and Anomaly Detection
To mitigate label dependency, recent work has increasingly explored unsupervised anomaly detection, in which models learn the distribution of nominal land cover and identify deviations without requiring explicit annotations [4]. Variational Autoencoders (VAEs) are frequently employed in this setting due to their ability to learn compact latent representations of complex imagery [36]. While many approaches rely on reconstruction error as an anomaly signal, prior studies have shown that reconstruction-based scores can be unreliable in fully unsupervised settings, particularly for high-capacity models [46]. More recent work has therefore emphasized the use of latent representations for comparing imagery across time, framing anomaly detection as a form of change detection in latent space [10]. This approach is particularly relevant in conflict zones, where unprecedented events and limited historical labels make representation-based change detection more robust than reconstruction-based anomaly scoring.
@@ 2.4 Sensor Selection and Data Processing
Satellite sensor choice determines the trade-off between spatial resolution and temporal frequency. Landsat and Sentinel-2 provide valuable multispectral data for burn scar analysis; however, their revisit intervals of 5–10 days often miss short-lived fire events [59]. Synthetic Aperture Radar (SAR) can penetrate cloud cover to assess post-fire land changes, though it lacks the spectral signatures provided by optical sensors [49]. PlanetScope imagery offers a balance, providing 3 m resolution and near-daily revisit cadences. However, leveraging this data for conflict monitoring requires robust preprocessing and computationally efficient models capable of handling large daily data volumes, factors that have limited its adoption in fully unsupervised, near–real-time fire detection workflows [12, 20, 8].
@@ 3 Study area
This study focuses on five conflict-related fire incidents in the El Fasher and Khartoum regions of Sudan. In Khartoum State, an airstrike reportedly carried out by the Sudanese Armed Forces (SAF) on May 25, 2024, targeted the Gandahar Market in Omdurman. The Conflict Observatory documented damage to at least 31 structures resulting from this incident [25]. PlanetScope imagery acquired on the day of the strike shows a visible smoke plume over the market area, consistent with active burning.
In North Darfur, political and security tensions escalated in late March 2024, with both the SAF and the Rapid Support Forces (RSF) preparing for confrontation in El Fasher by early May. According to the Conflict Observatory, at least 7,800 structures in eastern El Fasher were affected by fighting beginning on May 9, 2024, with NASA FIRMS data indicating widespread fire activity and extensive burning by May 11 [53]. Earlier, in April 2024, RSF-led forces attacked approximately 17 villages surrounding El Fasher. These attacks caused significant civilian casualties and widespread destruction of residential structures [57]. PlanetScope imagery captured active fires in several of these locations on April 13, 2024, including the villages of Jaranga, Muqrin, and Sarafaya [54].
Table 1: Details of satellite imagery used for the study area.
Place Area Incident Before-imagery After-imagery
($km^{2}$) date date date
Gandahar Market 4.37 May 25, 2024 May 13, 2024 May 25, 2024
El Fasher 5.41 May 09-11, 2024 May 02, 2024 May 11, 2024
Muqrin 4.52 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
Jaranga 2.31 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
Sarafaya 2.16 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
Table 1 summarizes the imagery and incident details for all case studies considered in this work. Additionally, Figure 2 illustrates the geographic distribution of the study areas across Sudan. These incidents were selected based on independently reported fire-related damage documented in [32, 57, 33], and manual evaluation and verification conducted by analysts associated with the Sudan Conflict Observatory project.
@@ 4.1 Model architecture
In this study, we adapted and retrained RaVAEn [56], a lightweight convolutional Variational Autoencoder (VAE) framework originally developed for unsupervised detection of disaster-related surface anomalies across multiple hazard types, including floods, fire burn scars, landslides, and hurricanes. Rather than being tailored to a specific phenomenon, RaVAEn learns a compact latent representation of nominal surface conditions and identifies anomalous patterns as deviations from this learned distribution, making it suitable for detecting previously unseen or irregular events such as conflict-related fires. The original model was designed for 10-band satellite imagery at 10 m spatial resolution. We retrained it for 3 m resolution imagery with four spectral channels (RGB and near-infrared) to better capture fine-scale spatial heterogeneity associated with small, fragmented fire events in urban conflict settings.
@@
To accommodate scale-variant spatial patterns, standard strided convolutions in the initial stages are replaced with multi-scale convolutional blocks analogous to Atrous Spatial Pyramid structures [40]. These blocks use parallel $3\times 3$ convolutions with increasing dilation rates, fused via a $1\times 1$ convolution. This design enables the encoder to capture multi-scale contextual dependencies and expand its effective receptive field while preserving spatial detail [17, 64]. Anti-aliased downsampling, or BlurPool [65], implemented as a stride-1 convolution followed by low-pass filtering and pooling, mitigates the loss of high-frequency information and improves shift equivariance when processing higher-resolution imagery. Each downsampling operation is followed by residual convolutional blocks with leaky ReLU activations and normalization layers, while deeper stages retain a conventional residual structure as higher-level semantic features are less sensitive to pixel-level resolution.
The final encoder feature map is aggregated using global pooling and projected onto a 128-dimensional latent space parameterized by a diagonal Gaussian distribution, regularized with a $\beta$-VAE objective ($\beta$=1). The decoder mirrors the encoder hierarchy using nearest-neighbor upsampling and convolutional layers, with a linear activation in the output layer. The model is trained end-to-end using the Adam optimizer with a learning rate of 0.001 and a batch size of 128, minimizing a combination of mean squared reconstruction loss and Kullback–Leibler divergence. To improve generalization across spatial resolutions, a scale-augmented training strategy is applied in which imagery is stochastically resampled to simulate continuous transitions in ground sampling distance [37]. This resampling is applied only during training; inference operates directly on native-resolution imagery, producing resolution-invariant feature embeddings.
While VAEs are commonly used for unsupervised anomaly detection by thresholding reconstruction error for individual inputs [5], prior work has shown that reconstruction error can be an unreliable indicator of anomalous behavior in fully unsupervised settings [46]. Motivated by these findings, we do not use reconstruction error as the anomaly signal. Instead, the VAE is used as a representation learning framework to embed surface conditions into a compact latent space, in which changes between temporally paired observations can be quantified for change detection.